# PepCompass: Navigating peptide embedding spaces using Riemannian Geometry

### v4

@@ Appendix D Proof of Theorem 1
Our manifolds of interest, $M_{z}^{\kappa}(\alpha),\ \alpha\in(0,1),\ z\in\mathcal{Z}$, are open subsets of $M^{\kappa}_{z}$ and therefore non-compact. To define Brownian motion in this setting, one needs to control the behaviour of paths near the boundary. Classical approaches include: (i) compactification [55], (ii) stopping the process at the boundary [28], or (iii) reflecting it [17]. In our work we adopt a compactification strategy via a smooth spherical cap (see Figure 6), followed by stopping on the boundary of a natural embedding of $M^{\kappa}_{z}(\alpha)$.
@@
Figure 6: Sphere-cap manifold compactification $\overline{M^{\kappa}_{z}}$. The figure illustrates how the $M_{z}^{\kappa}$ manifold is compactified by capping off the spherical domain with the appropriate boundary conditions.
@@
Let $\overline{M^{\kappa}_{z}(\alpha)}$ be the spherical-cap compactification of $M^{\kappa}_{z}(\alpha)$. Schwarz et al. [42] showed that the non-stopped version of Algorithm 1, extended to $\overline{M^{\kappa}_{z}(\alpha)}$, produces a process $\overline{Z^{\epsilon}}$ that converges to $B$ in the Skorokhod topology on $\overline{M^{\kappa}_{z}(\alpha)}$.
@@
Convergence in Skorokhod topology does not automatically imply convergence of stopping times (see Appendix D.1 for a counterexample). However, for the high-probability event
@@
Let $r(x)=d_{M_{z}^{\kappa}}(B_{0},x)$ and write $r_{t}:=r(B_{t})$. On $M_{z}^{\kappa}$ the function $r$ is smooth. The semimartingale decomposition of $r_{t}$ (see, e.g., [27, Eq. (3.6.1)]) gives
@@
On the event $\{T_{R_{z}^{\kappa}}\leq T\}$ we have $t=T_{R_{z}^{\kappa}}$ and $r_{t}\geq R_{z}^{\kappa}$, hence from equation 14
@@
The lemma shows that the probability of exiting the ball of radius $R^{\kappa}_{z}$ before time $T$ decays exponentially in $\tfrac{(R^{\kappa}_{z})^{2}}{T}$, up to curvature- and rank-dependent constants. Intuitively, this means that with overwhelming probability the Riemannian Brownian motion (and hence our random walk in the $\epsilon\to 0$ limit) remains confined inside $B_{{M^{\kappa}_{z}}(}(z,R^{\kappa}_{z})$ for all $t\leq T$. This high-probability control is what allows us to restrict attention to the event $C^{T}_{\kappa,z}$ in the proof of Theorem 1.
D.1 On stopping times and Skorohod convergence
An important subtlety in the proof of Theorem 1 is that convergence of processes in the Skorohod topology does not, in general, imply convergence of associated stopping times. Figure 7 illustrates this phenomenon with a simple deterministic example.
@@ D.1 On stopping times and Skorohod convergence
In the proof of Theorem 1 we avoid this issue by restricting to the high-probability set $C^{T}_{\kappa,z}$ where the Brownian path remains in the interior of the ball. On this event the stopping times agree with $T$, ensuring consistency with the limiting process.
Appendix E SORBES implementation details
@@
We adopt the efficient implementation of Schwarz et al. [42], which exploits a thin SVD of the decoder Jacobian to orthogonalize tangent directions.
@@
Computing Christoffel symbols directly requires evaluating first derivatives of the metric, which is computationally expensive and numerically unstable. Instead, we use an extrinsic approach: the covariant derivative of a curve can be obtained from its Euclidean acceleration in the ambient space, projected back onto the tangent space [16]. Concretely, for a point $z\in\mathcal{Z}$, $z^{\prime}\in W^{\kappa}_{z}$ and a probe radius $\rho>0$, we approximate the extrinsic acceleration of a decoded curve along $v\in T_{z^{\prime}}M_{z}^{\kappa}$ by a second-order central difference:
@@
where $a\vee b=\max\{a,b\}$. This guarantees that the step size is never smaller than the nominal $\epsilon$, but shrinks adaptively whenever the second-order correction is large.
@@
We refer to the resulting algorithm with adaptive step size control and extrinsic approximation of Christoffel symbols (Sec. 2.3) as SORBES-SE. This variant is numerically stable in ill-conditioned regions of the decoder geometry while preserving the efficiency of the original scheme.
@@
21:end while
22:return $(z^{\epsilon}_{i})_{0\leq i\leq\texttt{step}},\ \sigma$
Appendix F MUTANG - Mutation Enumeration in Tangent Space
@@
18:return $\mathcal{C}(\mathtt{p}(z))$
Appendix G PoGS
@@
4:  Compute energy $\mathcal{E}_{\lambda}(Z)$ as in equation 6
@@
6:end for
7:return $\{\mathtt{p}(z_{k})\}_{k=0}^{N}$
Appendix H APEX-potential for PoGS
The APEX predictor [53] estimates minimum inhibitory concentration (MIC) values against $11$ bacterial strains, but it operates on concrete peptide sequences. In Potential-Minimizing Geodesic Search (PoGS), optimization proceeds over latent-space chords, i.e., intermediate points $z$ that decode to position-factorized distributions over peptides rather than single sequences:
@@ Appendix H APEX-potential for PoGS
Peptides from the HydrAMP training set  [50] were encoded into latent codes $z$. In order to obtain multiple distributions from a single peptide, we then created four clones $z^{\prime}$ of latent codes $z$. We applied a 2×2 perturbation scheme: two clones were injected with Gaussian noise $N(0,0.05)$, and two were left unchanged. Finally, these four latent codes were decoded to $\operatorname{Dec}(z^{\prime})$, using a softmax scaling with temperature of 1.0 for one pair (noisy and non-noisy) and a temperature of 1.5 to the other pair, resulting in four distributions per peptide. This yielded 1,060,000 peptide distributions in total. For each $\operatorname{Dec}(z^{\prime})$, we enumerated the $N=20$ most-probable sequences
@@
Restricting to the three target E. coli strains (index set $\mathcal{I}_{\textit{E.~coli}}$), the scalar property potential used by PoGS is
$\Phi(X)=\mathbf{1}^{\top}\big[\widehat{m}^{\mathrm{z}}(X)\big]_{\mathcal{I}_{\textit{E.~coli}}}\,.$
Appendix I LE-BO Hyperparameters
@@
    • Following the approach of [18], as a surrogate model, we use a Gaussian Process $\operatorname{GP}$ with the Tanimoto similarity kernel [47], applied to the MAP4 fingerprints of peptides [11].
@@
    • Mutation threshold $\theta_{\text{mut}}$
@@
Appendix J Wet-lab validation
J.1 Peptide synthesis and characterization
@@
J.2 Bacterial Strains and Growth Conditions
The bacterial panel utilized in this study consisted of the following pathogenic strains: Acinetobacter baumannii ATCC 19606; A. baumannii ATCC BAA-1605 (resistant to ceftazidime, gentamicin, ticarcillin, piperacillin, aztreonam, cefepime, ciprofloxacin, imipenem, and meropenem); Escherichia coli ATCC 11775; E. coli AIC221 [MG1655 phnE_2::FRT, polymyxin-sensitive control]; E. coli AIC222 [MG1655 pmrA53 phnE_2::FRT, polymyxin-resistant]; E. coli ATCC BAA-3170 (resistant to colistin and polymyxin B); Enterobacter cloacae ATCC 13047; Klebsiella pneumoniae ATCC 13883; K. pneumoniae ATCC BAA-2342 (resistant to ertapenem and imipenem); Pseudomonas aeruginosa PAO1; P. aeruginosa PA14; P. aeruginosa ATCC BAA-3197 (resistant to fluoroquinolones, $\beta$-lactams, and carbapenems); Salmonella enterica ATCC 9150; S. enterica subsp. enterica Typhimurium ATCC 700720; Bacillus subtilis ATCC 23857; Staphylococcus aureus ATCC 12600; S. aureus ATCC BAA-1556 (methicillin-resistant); Enterococcus faecalis ATCC 700802 (vancomycin-resistant); and Enterococcus faecium ATCC 700221 (vancomycin-resistant). P. aeruginosa strains were propagated on Pseudomonas Isolation Agar, whereas all other species were maintained on Luria-Bertani (LB) agar and broth. For each assay, cultures were initiated from single colonies, incubated overnight at 37 °C, and subsequently diluted 1:100 into fresh medium to obtain cells in mid-logarithmic phase.
J.3 Minimal Inhibitory Concentration (MIC) determination
@@
J.4 Detailed wet-lab validation results

### v5

@@ D.3 Extension to biologically meaningful ambient metrics by mutation kernels
Our manifolds of interest, $M_{z}^{\kappa}(\alpha),\ \alpha\in(0,1),\ z\in\mathcal{Z}$, are open subsets of $M^{\kappa}_{z}$ and therefore non-compact. To define Brownian motion in this setting, one needs to control the behaviour of paths near the boundary. Classical approaches include: (i) compactification (Wang, 2011), (ii) stopping the process at the boundary (Hsu, 2002), or (iii) reflecting it (Du and Hsu, 2021). In our work we adopt a compactification strategy via a smooth spherical cap (see Figure 6), followed by stopping on the boundary of a natural embedding of $M^{\kappa}_{z}(\alpha)$.
@@
Figure 6: Sphere-cap manifold compactification $\overline{M^{\kappa}_{z}}$. Illustration of how the $M_{z}^{\kappa}$ manifold is compactified by capping off the spherical domain with the appropriate boundary conditions.
@@
Let $\overline{M^{\kappa}_{z}(\alpha)}$ be the spherical-cap compactification of $M^{\kappa}_{z}(\alpha)$. Schwarz et al. (2023) showed that the non-stopped version of Algorithm 1, extended to $\overline{M^{\kappa}_{z}(\alpha)}$, produces a process $\overline{Z^{\epsilon}}$ that converges to $B$ in the Skorokhod topology on $\overline{M^{\kappa}_{z}(\alpha)}$.
@@
Convergence in Skorokhod topology does not automatically imply convergence of stopping times (see Appendix E.1 for a counterexample). However, for the high-probability event
@@
Let $r(x)=d_{M_{z}^{\kappa}}(B_{0},x)$ and write $r_{t}:=r(B_{t})$. On $M_{z}^{\kappa}$ the function $r$ is smooth. The semimartingale decomposition of $r_{t}$ (see, e.g., Eq. 3.6.1 in (Hsu, 2002)) gives
@@
On the event $\{T_{R_{z}^{\kappa}}\leq T\}$ we have $t=T_{R_{z}^{\kappa}}$ and $r_{t}\geq R_{z}^{\kappa}$, hence from (14)
@@
The lemma shows that the probability of exiting the ball of radius $R^{\kappa}_{z}$ before time $T$ decays exponentially in $\tfrac{(R^{\kappa}_{z})^{2}}{T}$, up to curvature- and rank-dependent constants. Intuitively, this means that with overwhelming probability the Riemannian Brownian motion (and hence our random walk in the $\epsilon\to 0$ limit) remains confined inside $B_{{M^{\kappa}_{z}}(}(z,R^{\kappa}_{z})$ for all $t\leq T$. This high-probability control is what allows us to restrict attention to the event $C^{T}_{\kappa,z}$ in the proof of Theorem 2.1.
E.1 On stopping times and Skorohod convergence
An important subtlety in the proof of Theorem 2.1 is that convergence of processes in the Skorohod topology does not, in general, imply convergence of associated stopping times. Figure 7 illustrates this phenomenon with a simple deterministic example.
@@ E.1 On stopping times and Skorohod convergence
In the proof of Theorem 2.1 we avoid this issue by restricting to the high-probability set $C^{T}_{\kappa,z}$ where the Brownian path remains in the interior of the ball. On this event the stopping times agree with $T$, ensuring consistency with the limiting process.
Appendix F SORBES implementation details
@@
We adopt the efficient implementation of Schwarz et al. (2023), which exploits a thin SVD of the decoder Jacobian to orthogonalize tangent directions.
@@
Computing Christoffel symbols directly requires evaluating first derivatives of the metric, which is computationally expensive and numerically unstable. Instead, we use an extrinsic approach: the covariant derivative of a curve can be obtained from its Euclidean acceleration in the ambient space, projected back onto the tangent space (Flaherty and do Carmo, 2013). Concretely, for a point $z\in\mathcal{Z}$, $z^{\prime}\in W^{\kappa}_{z}$ and a probe radius $\rho>0$, we approximate the extrinsic acceleration of a decoded curve along $v\in T_{z^{\prime}}M_{z}^{\kappa}$ by a second-order central difference:
@@
where $a\vee b=\max\{a,b\}$. This guarantees that the step size is never greater than the nominal $\epsilon$, but shrinks adaptively whenever the second-order correction is large.
@@
We refer to the resulting algorithm with adaptive step size control and extrinsic approximation of Christoffel symbols (Section 2.3) as SORBES-SE. This variant is numerically stable in ill-conditioned regions of the decoder geometry while preserving the efficiency of the original scheme.
@@
21:end whilereturn $(z^{\epsilon}_{i})_{0\leq i\leq\texttt{step}},\ \sigma$
Appendix G MUTANG - Mutation Enumeration in Tangent Space
@@
Appendix H Computational Scalability of Local Enumeration
Because SORBES-SE uses a finite-difference approximation, we can estimate the inference-time cost of Local Enumeration as a function of peptide length and latent dimensionality. Let $t_{\text{decoder}}$ denote the decoder forward-pass time, $k$ the latent dimension, $n$ the maximum peptide length, and $a$ the alphabet size. The pessimistic computational cost of the Local Enumeration (LE) step can then be approximated as follows:
  • Finite-difference evaluations for the Jacobian ($k$ evaluations) and the Christoffel symbols (2 evaluations): $(k+2),t_{\text{decoder}}$.
  • Singular value decomposition (SVD): $\mathcal{O}(kna)$.
  • Projection of the extrinsic acceleration (second-order correction): $\mathcal{O}(kna)$.
  • MUTANG computation: $\mathcal{O}(kn)$.
Overall, a Local Enumeration step scales linearly with the latent dimension $k$, peptide length $n$, and decoder evaluation time $t_{\text{decoder}}$.
@@ Appendix H Computational Scalability of Local Enumeration
Appendix I PoGS
@@
4:  Compute energy $\mathcal{E}_{\lambda}(Z)$ as in (6)
@@
6:end forreturn $\{\mathtt{p}(z_{k})\}_{k=0}^{N}$
Appendix J APEX-potential for PoGS
The APEX predictor (Wan et al., 2024) estimates minimum inhibitory concentration (MIC) values against $11$ bacterial strains, but it operates on concrete peptide sequences. In Potential-Minimizing Geodesic Search (PoGS), optimization proceeds over latent-space chords, i.e., intermediate points $z$ that decode to position-factorized distributions over peptides rather than single sequences:
@@
Peptides from the HydrAMP training set (Szymczak et al., 2023) were encoded into latent codes $z$. In order to obtain multiple distributions from a single peptide, we then created four clones $z^{\prime}$ of latent codes $z$. We applied a 2×2 perturbation scheme: two clones were injected with Gaussian noise $N(0,0.05)$, and two were left unchanged. Finally, these four latent codes were decoded to $\operatorname{Dec}(z^{\prime})$, using a softmax scaling with temperature of 1.0 for one pair (noisy and non-noisy) and a temperature of 1.5 to the other pair, resulting in four distributions per peptide. This yielded 1,060,000 peptide distributions in total. For each $\operatorname{Dec}(z^{\prime})$, we enumerated the $N=20$ most-probable sequences
@@
Restricting to the three target E. coli strains (index set $\mathcal{I}_{\textit{E.\penalty 10000\ coli}}$), the scalar property potential used by PoGS is
$\Phi(X)=\mathbf{1}^{\top}\big[\widehat{m}^{\mathrm{z}}(X)\big]_{\mathcal{I}_{\textit{E.\penalty 10000\ coli}}}\,.$
Appendix K PoGS ablation study
K.1 Analysis of results for different choice of metric and prototypes set
To assess sensitivity of PoGS to the choice of the metric as well as to the choice of prototypes, we compared our implementation with two alternatives: using amino-acid probabilities instead of logits, and straight (Euclidean) interpolation, on a new set of 60 prototypes (Table 3). Logits yielded better performance in terms of potential, counts of seeds and wells, supporting our choice. Importantly, both metrics outperformed straight interpolation, and PoGS achieved even better results on these new prototypes than those reported in Table 1. These results show that PoGS improves results regardless of the metric used or the prototype set.
Table 3: PoGS with the original metric compared to PoGS with a metric on decoded probabilities, as well as to straight interpolation, for a different set of prototypes than in Table 1.
Method Latent Length Ambient Length (orig.) Ambient Length (probs) Peptide Path Length Potential Seeds Wells
Straight interpolation $6.50\pm 1.34$ $3448.7\pm 403.8$ $4.76\pm 4.94$ $47.0\pm 20.9$ $732.8\pm 272.7 11$$17\pm 42$ $41\pm 14$
PoGS (with probs metric) $11.48\pm 3.36$ $17227.93\pm 147.8$ $1.67\pm 1.16$ $42.7\pm 18.1$ $1729.9\pm 271.4$ $176\pm 43$ $67\pm 14$
PoGS (original) $14.12\pm 3.07$ $5001.4\pm 236.7$ $9.97\pm 4.96$ $63.3\pm 29.3$ $2329.0\pm 420.0$ $188\pm 29$ $80\pm 20$
K.2 Analysis of results for different choice of potential
To evaluate PoGS’s sensitivity to the choice of potential and its applicability to multi-objective settings, we applied it to jointly maximize two physicochemical properties by assigning weight alphaalpha to the hydrophobicity and 1 – $\alpha$ to charge (Table 4). Across all values of $\alpha$, PoGS outperformed straight-line (Euclidean) interpolation. For this experiment, we selected a new set of 60 prototypes from the GRAMPA and DRAMP datasets, demonstrating that PoGS achieves superior performance regardless of both the potential used and the prototype set.
Method Latent Length Ambient Length Peptide Path Length Max Hydrophobicity Max Charge Maximized Multiobjective Potential
PoGS w. $\alpha=0.9$ $12.07\pm 2.57$ $4573.57\pm 19$ $66.86\pm 24.14$ $8.40\pm 0.72$ $15.81\pm 2.31$ $24.21\pm 2.42$
PoGS w. $\alpha=0.5$ $10.04\pm 2.05$ $3624.69\pm 185.09$ $57.66\pm 27.08$ $7.64\pm 0.63$ $15.75\pm 2.15$ $23.39\pm 2.24$
PoGS w. $\alpha=0.1$ $10.61\pm 2.13$ $3651.19\pm 190.39$ $55.93\pm 25.80$ $7.91\pm 0.64$ $14.76\pm 2.26$ $22.67\pm 2.35$
Straight line (Euclidean) $6.51\pm 1.34$ $3448.70\pm 403.80$ $47.00\pm 20.90$ $7.90\pm 0.64$ $13.76\pm 2.22$ $21.66\pm 2.35$
Table 4: Comparison of PoGS multiobjective optimization across different $\alpha$ values and Euclidean straight-line baselines.
Appendix L LE-BO Hyperparameters
@@
    • Following the approach of (Eberhardt et al., 2024), as a surrogate model, we use a Gaussian Process $\operatorname{GP}$ with the Tanimoto similarity kernel (Szedmak and Bach, 2020), applied to the MAP4 fingerprints of peptides (Capecchi et al., 2020).
@@
    • Mutation threshold $\theta_{\text{mut}}=10^{-6}$.
@@
Appendix M LE-BO time and memory profiling
To quantify the computational gains of our method, we measured the average runtime of LE-BO over 10 iterations and 6 seeds, and compared it to SAASBO ((Eriksson and Jankowiak, 2021)) under the same conditions (Table 5). LE-BO required 8× less time per iteration and used 1.5× less memory. Moreover, Local Enumeration accounted for only 35% of the total runtime, underscoring its efficiency. Average execution time of a single optimization run of LE-Bo with 1400 iterations was 1h20m (±30m). All computations were performed for the HydrAMP model with dimension 64 and measured on a Mac Mini M4Pro machine with 24GB of RAM.
Method Local Enumeration Time / Iteration (s) Total Iteration Time (s) Memory (MB)
LE-BO $1.03\pm 0.31$ $2.89\pm 2.21$ $1490\pm 210$
SAASBO ((Eriksson and Jankowiak, 2021)) N/A $3.31\pm 9.94 22$$248\pm 12$
Table 5: Runtime and memory comparison between LE-BO and SAASBO ((Eriksson and Jankowiak, 2021)).
Appendix N LE-BO Ablation study
N.1 Analysis of results for different $\kappa_{SORBES}$ and $\kappa_{MUTANG}$
Additional ablation study show that alternative $\kappa$ values can even outperform those originally selected, demonstrating that LE-BO’s performance can be further enhanced through targeted hyperparameter tuning  6. An ablation with respect to $\alpha$ is unnecessary because, with only a single SORBES-SE step, $\alpha$ has no effect on the search process.
$\boldsymbol{\kappa_{\text{mutang}}}$ $\boldsymbol{\kappa_{\text{sorbes}}}$ FL14 KY14 KF16 KK16 mammuthusin-3 hydrodamin-2 Avg. Diff. from LE-BO
$10^{-8}$ $10^{-3}$ $0.74\pm 0.28$ $0.83\pm 0.44$ $0.56\pm 0.29$ $0.58\pm 0.30$ 0.44 $\pm$ 0.29 0.38 $\pm$ 0.13 $0.04\pm 0.18$
$10^{-4}$ $10^{-3}$ $0.53\pm 0.26$ $5\pm 0.39 0.66±0$$6\pm 0.11$ 0.36 $\pm$ 0.09 $0.60\pm 0.69$ $45\pm 0.14 -0.0$1 $\pm$ 0.12
$10^{-8}$ $10^{-1}$ 0.41 $\pm$ 0.11 $0.65\pm 0.44$ 0.38 $\pm$ 0.14 $0.73\pm 0.34$ $0.58\pm 0.21$ $45\pm 0.11 0.0$$01\pm 0.19$
$10^{-4}$ $10^{-1}$ $0.83\pm 0.20$ $0.54\pm 0.31$ $0.72\pm 0.29$ $0.66\pm 0.31$ $0.60\pm 0.45$ $0.63\pm 0.40$ $0.12\pm 0.07$
$10^{-6}$ (orig.) $10^{-2}$ (orig.) $0.604\pm 0.22$ 0.502 $\pm$ 0.24 $0.600\pm 0.29$ $0.498\pm 0.14$ $0.498\pm 0.38$ $581\pm 0.34 0.0$$0\pm 0.0$
Table 6: LE-BO performance for different hyperparameter values. Best values of minimized MIC as predicted by APEX are bolded. The last column reports the average difference between the alternative hyperparameter setting and the original used in the manuscript.
N.2 Analysis of results of LE-BO without ROBOT ((Maus et al., 2023))
Ablation analysis of the model without ROBOT (Maus et al., 2023) confirms that ROBOT improves LE-BO for 4 out of 6 seeds (with average difference in log MIC of -0.03).
Method FL14 KY14 KF16 KK16 mammuthusin-3 hydrodamin-2 Avg. diff. from LE-BO
LE-BO w/o ROBOT $0.75\pm 0.41$ $0.52\pm 0.17$ 0.56 $\pm$ 0.11 $0.63\pm 0.28$ $0.57\pm 0.61$ 0.42 $\pm$ 0.20 $0.03\pm 0.12$
LE-BO 0.604 $\pm$ 0.22 0.502 $\pm$ 0.24 0.600 $\pm$ 0.29 0.498 $\pm$ 0.14 0.498 $\pm$ 0.38 0.581 $\pm$ 0.34 0.000 $\pm$ 0.000
Table 7: Comparison of LE-BO with and without ROBOT across six peptides and mean deviation from LE-BO baseline.
N.3 Analysis of results for other seeds
To show consistent gains compared to ablations across different seeds, we additionally performed an LE-BO ablation study on three further seeds (LL13, RC16, and KI21) from the PoGS optimization (Table 8). The results confirm that clear advantage of enabling mutations and of using SORBES-SE as the random walk method can be observed regardless of the chosen seeds. Taken together, all three LE-BO components - SORBES-SE, enabling mutation (MUTANG), and ROBOT - contribute meaningful improvements. However, enabling mutation with MUTANG provides the most substantial benefit: across all ablations, the LE-BO variants incorporating MUTANG consistently outperform their counterparts without it.
Walk Mutation LL13 RC16 KI21 Difference from LE-BO
Euclidean $\times$ $0.787\pm 0.408$ $141\pm 0.298 1.1$$170\pm 0.283$ $1.033\pm 0.174$
SORBES-SE $\times$ $009\pm 0.408 1.1$$28\pm 0.298 1.1$$120\pm 0.283$ $0.709\pm 0.204$
$\checkmark$ $1.343\pm 0.356$ $1.225\pm 0.275$ $0.803\pm 0.238$ $0.747\pm 0.125$
Euclidean $\checkmark$ $0.630\pm 0.190$ $0.751\pm 0.291$ $0.505\pm 0.199$ $0.253\pm 0.091$
LE-BO SORBES-SE $\checkmark$ 0.482 $\pm$ 0.226 0.458 $\pm$ 0.243 0.190 $\pm$ 0.191 0.000 $\pm$ 0.000
Table 8: Confirmation of improved performance of LE-BO compared to ablations for different seeds than in Table 2
N.4 Analysis of results for different oracles
To evaluate robustness across different oracles and to verify that additional peptide properties benefit from geometry-aware exploration, we successfully applied LE-BO and show that geometry-aware exploration improves prototype peptides for three tasks:
  1. 1.
    minimize MIC using DEEP-AMP (Pandi et al., 2023) regressor other than APEX as oracle,
  2. 2.
    minimize toxicity using ToxiPrep (Guan et al., 2025) classification probabilities as oracle,
  3. 3.
    maximize hydrophobicity computed in the Eisenberg scale (Eisenberg, David and Weiss, Robert M. and Terwilliger, Thomas C. and Wilcox, William, 1982) as oracle.
KY14 KF16 KK16 FL14 mammuthusin-3 hydrodamin-2
log2(MIC) (seed) $2.00$ $5.06$ $1.85$ $1.82$ $4.02$ $5.09$
Euclidean LE-BO -0.66 $\pm$ 0.50 0.04 $\pm$ 0.19 -0.56 $\pm$ 0.65 -0.55 $\pm$ 0.46 -0.48 $\pm$ 0.72 -0.15 $\pm$ 0.56
LE-BO -0.80 $\pm$ 1.06 -0.47 $\pm$ 0.38 -1.20 $\pm$ 1.23 -2.35 $\pm$ 0.41 -0.77 $\pm$ 0.75 -0.93 $\pm$ 0.89
Table 9: Minimization of $\text{log}_{2}$(MIC) values using LE-BO with DEEP-AMP (Pandi et al., 2023) as oracle.
KY14 KF16 KK16 FL14 mammuthusin-3 hydrodamin-2
Toxicity (seed) $0.8789$ $0.9283$ $0.9044$ $0.9932$ $0.5193$ $0.0522$
Euclidean LE-BO 0.014 $\pm$ 0.002 0.016 $\pm$ 0.003 0.016 $\pm$ 0.005 0.015 $\pm$ 0.001 0.014 $\pm$ 0.002 0.016 $\pm$ 0.002
LE-BO 0.0115 $\pm$ 0.0007 0.0117 $\pm$ 0.0012 0.0120 $\pm$ 0.0022 0.0122 $\pm$ 0.0015 0.0126 $\pm$ 0.0013 0.0135 $\pm$ 0.0012
Table 10: Minimization of toxicity values using LE-BO with toxicity probabilities returned by ToxiPep (Guan et al., 2025) as oracle.
KY14 KF16 KK16 FL14 mammuthusin-3 hydrodamin-2
Hydrophobicity (seed) 0.34-0.34 0.60-0.60 $0.03$ $0.12$ $0.17$ 0.33-0.33
Euclidean LE-BO 1.114 $\pm$ 0.091 1.208 $\pm$ 0.048 1.162 $\pm$ 0.112 1.193 $\pm$ 0.029 1.323 $\pm$ 0.031 0.963 $\pm$ 0.084
LE-BO 1.255 $\pm$ 0.013 1.270 $\pm$ 0.048 1.242 $\pm$ 0.067 1.265 $\pm$ 0.026 1.325 $\pm$ 0.022 0.953 $\pm$ 0.095
Table 11: Maximization of hydrophobicity using LE-BO with hydrophobicity values computed in Eisenberg scale (Eisenberg, David and Weiss, Robert M. and Terwilliger, Thomas C. and Wilcox, William, 1982) oracle.
Appendix O Geometrical Analysis of LE-BO
Across both HydrAMP and PepCVAE, the effective rank of the latent space is consistently much lower than the nominal latent dimension (Figure 4A,B). Furthermore, the effective rank exhibits a strong positive correlation with peptide length (Figure 4C,D; Appendix B), indicating increasing geometric complexity for longer peptides.
Figure 8 analyzes how this latent dimensionality affects local candidate enumeration. LE-BO with SORBES-SE and mutation enabled (Figure 8A,E) identifies substantially more candidates in local enumeration as the $\kappa$-stable dimension increases, corresponding to longer peptides. In contrast, Euclidean-based search exhibits the opposite trend, producing fewer candidates in higher-dimensional regions (Figure 8B,C,F,G). Since the proposed optimization procedure relies on dense local peptide neighborhoods, this behavior explains the weaker empirical performance of Euclidean approaches.
Quantitatively, LE-BO with SORBES-SE and mutation generates significantly more candidates on average (16,730 $\pm$ 14,786; Figure 8A,E) than Euclidean search with mutation (442 $\pm$ 151; Figure 8B,F). Even without mutation, SORBES-SE yields higher candidate counts (17 $\pm$ 5; Figure 8D,H) than its Euclidean counterpart (13 $\pm$ 5; Figure 8C,G). Finally, the $\kappa$-stable latent dimension is strongly negatively correlated with peptide charge (Spearman r=0.55r=-0.55) and aromaticity (r=0.57r=-0.57; Figure 9). This suggests that peptides associated with higher latent dimensionality tend to exhibit higher solubility but reduced membrane affinity and antimicrobial activity. By enumerating substantially more candidates in these high-dimensional regions, LE-BO increases the likelihood that such unfavorable properties can be corrected during optimization.
Figure 8: Relationships between Local Enumeration (LE) result size and sequence properties under different ablations of LE-BO walk and mutation strategies. Each panel (A–H) shows scatter plots comparing LE size with either peptide length (top row) or $\kappa$-stable dimension (bottom row) for four experimental LE-BO conditions: SORBES-SE walk with mutation (default LE-BO implementation) (A, E), Euclidean walk with mutation (B, F), Euclidean walk without mutation (C, G), and SORBES-SE walk without mutation (D, H). Titles report the mean ± standard deviation of the LE size for each condition. Together, these comparisons highlight how walk geometry and mutation choice influence the size and structure of local neighborhoods explored during Local Enumeration.
Figure 9: Relationship between peptide physicochemical properties and latent dimension. (A) Aromaticity versus embedding dimension, showing that peptides with lower aromatic residue content tend to occupy higher-dimensional regions of the embedding space. (B) Net charge at pH 7.4 versus embedding dimension, highlighting a similar trend in which peptides with lower cationic charge are more common at higher dimensions.
Appendix P Wet-lab validation
P.1 Experimental setup
P.2 Peptide synthesis and characterization
@@
P.3 Bacterial Strains and Growth Conditions
The bacterial panel utilized in this study consisted of the following pathogenic strains (see Table 12: Acinetobacter baumannii ATCC 19606; A. baumannii ATCC BAA-1605 (resistant to ceftazidime, gentamicin, ticarcillin, piperacillin, aztreonam, cefepime, ciprofloxacin, imipenem, and meropenem); Escherichia coli ATCC 11775; E. coli AIC221 [MG1655 phnE_2::FRT, polymyxin-sensitive control]; E. coli AIC222 [MG1655 pmrA53 phnE_2::FRT, polymyxin-resistant]; E. coli ATCC BAA-3170 (resistant to colistin and polymyxin B); Enterobacter cloacae ATCC 13047; Klebsiella pneumoniae ATCC 13883; K. pneumoniae ATCC BAA-2342 (resistant to ertapenem and imipenem); Pseudomonas aeruginosa PAO1; P. aeruginosa PA14; P. aeruginosa ATCC BAA-3197 (resistant to fluoroquinolones, $\beta$-lactams, and carbapenems); Salmonella enterica ATCC 9150; S. enterica subsp. enterica Typhimurium ATCC 700720; Bacillus subtilis ATCC 23857; Staphylococcus aureus ATCC 12600; S. aureus ATCC BAA-1556 (methicillin-resistant); Enterococcus faecalis ATCC 700802 (vancomycin-resistant); and Enterococcus faecium ATCC 700221 (vancomycin-resistant). P. aeruginosa strains were propagated on Pseudomonas Isolation Agar, whereas all other species were maintained on Luria-Bertani (LB) agar and broth. For each assay, cultures were initiated from single colonies, incubated overnight at 37 °C, and subsequently diluted 1:100 into fresh medium to obtain cells in mid-logarithmic phase.
Table 12: Bacterial strains used for experimental validation of antimicrobial peptide libraries. Strains marked with MDR are multidrug-resistant clinical isolates.
ID Bacterial Strain
AB1 A. baumannii ATCC 19606
@@
EC1 E. cloacae ATCC 13047
EC2 E. coli ATCC 11775
EC3 E. coli AIC221